factor completely
150r^2+360rg+216g^2

Found 2 solutions by stanbon, fractalier:
Answer by stanbon(75887)

factor completely
150r^2+360rg+216g^2
-------
= 2[75r^2 + 180rg + 108q^2]
----
= 2*3[25r^2 + 60rg + 36]
----
= 6(5r+6)(5r+6)
==================
Cheers,
Stan H.
==================

Answer by fractalier(6550)

From 150r^2+360rg+216g^2,
we first pull out the greatest common factor, which is 6, and get
6(25r^2 + 60rg + 36g^2)
What remains is a perfect square trinomial...
= 6(5r + 6g)^2